# GP-GS: Gaussian Processes for Enhanced Gaussian Splatting

### v2

@@ Abstract
3D Gaussian Splatting has emerged as an efficient photorealistic novel view synthesis method. However, its reliance on sparse Structure-from-Motion (SfM) point clouds consistently compromises the scene reconstruction quality. To address these limitations, this paper proposes a novel 3D reconstruction framework Gaussian Processes Gaussian Splatting (GP-GS), where a multi-output Gaussian Process model is developed to achieve adaptive and uncertainty-guided densification of sparse SfM point clouds. Specifically, we propose a dynamic sampling and filtering pipeline that adaptively expands the SfM point clouds by leveraging GP-based predictions to infer new candidate points from the input 2D pixels and depth maps. The pipeline utilizes uncertainty estimates to guide the pruning of high-variance predictions, ensuring geometric consistency and enabling the generation of dense point clouds. The densified point clouds provide high-quality initial 3D Gaussians to enhance reconstruction performance. Extensive experiments conducted on synthetic and real-world datasets across various scales validate the effectiveness and practicality of the proposed framework.
@@ Introduction
Related Work
Gaussian Processes
@@
Methodology
Overview
This paper proposes a novel framework GP-GS that enhances 3D Gaussians initialization, thereby improving 3DGS rendering quality. The overview of GP-GS is presented in Figure 2. First, addressing the limitations of sparse SfM point clouds, we propose the MOGP model to formulate point cloud densification as a regression problem. (Sec. Multi Output Gaussian Process). The training strategy of MOGP is introduced in Appendix A. Next, we adaptively densify sparse SfM point clouds and propose an uncertainty-based filtering strategy to remove high-uncertainty predictions. (Sec. Point Cloud Densification).
Multi Output Gaussian Process
@@
where $\Gamma(\cdot)$ is the Gamma function, $\sigma^{2}$ is the variance, and $K_{\nu}(\cdot)$ is a modified Bessel function of the second kind. The parameter $\nu$ modulates how smooth the function can be: smaller $\nu$ values permit abrupt transitions, while larger $\nu$ values enforce smoother spatial changes. Details of tuning $\nu$ on various datasets can be found in Table 5. Our MOGP training procedure can be found in Appendix A Algorithm 1.
Point Cloud Densification
The process of point cloud densification involves an adaptive neighbourhood-based sampling strategy, which selectively identifies pixels as candidate inputs for the MOGP model (in section Multi Output Gaussian Process). The MOGP model subsequently predicts additional point clouds and applies uncertainty filtering to enhance the density.
@@
An overview of the uncertainty-based filtering procedure can be found in Appendix A Algorithm 2.
Experiments
Dataset and Implementation Details
@@
Results of Gaussian Processes
@@
Quantitative Results To ensure the best performance of MOGP for densification, we evaluated MOGP model with different hyperparameters on our Pixel-to-Point-Cloud dataset, e.g., $\nu$ with values such as ($\nu\in\{0.5,1.5,2.5\}$). The data were split into an 80% training subset and a 20% testing subset, and the model was subsequently assessed using $R^{2}$, RMSE, and CD as we mentioned in section Dataset and Implementation Details. The average outcomes of this analysis are summarized in Table 1, which shows that $\nu=0.5$ consistently yields the most favourable results. Consequently, we adopt $\nu=0.5$ as our default configuration for subsequent experiments. Detailed metrics across multiple datasets for this kernel selection procedure are presented in Appendix Results of Gaussian Processes Table 5.
@@
To validate the effectiveness of our variance-based filtering algorithm, we analyze the variance distributions across multiple Pixel-to-Point datasets, including NeRF Synthetic, Tanks & Temples, and Mip-NeRF 360, due to their diverse characteristics. The reduction in variance in all datasets suggests that our filtering method generalizes well to different types of 3D scenes, improving overall point cloud quality and enabling more accurate scene representation, as shown in Figure 5, the left column shows the spatial and colour variances before filtering on Tanks & Temples (Truck), the right column presents the variances after applying our filter. The variances in spatial components $(x,y,z)$ and colour channels $(r,g,b)$ are significantly reduced, demonstrating that our filtering process effectively removes high-variance noisy points while preserving scene consistency. More results on various datasets can be found in Appendix A.
@@
Quantitative and Qualitative Results on NVS
@@
Qualitative Results As shown in Figure 6, we compare our method with 3DGS* on datasets Tanks & Temples (Truck), Mip-NeRF 360 (Flowers and Bicycles), and NeRF Synthetic (Lego), we provide additional results in Appendix A for each scene to further substantiate our conclusions. Our method demonstrates superior performance in NVS by effectively reconstructing fine-grained details, reducing artifacts, and enhancing occlusion handling. The visual comparisons highlight key challenges in 3DGS* rendering, including occlusions, high-frequency textures (e.g., object edges and surface details), and lighting variations. Notably, the zoom-in regions (marked with red and blue boxes) illustrate our model’s ability to recover fine structures more accurately than existing approaches. These results further validate the robustness of our method in handling complex scenes, offering improved perceptual quality and structural consistency across diverse datasets.
Conclusion
@@ Appendix A Appendix
MOGP Training and Dynamic Filtering Procedure
@@
MOGP Training Loss
@@
More MOGP Results
@@
More Novel View Synthesis Results

### v3

@@ Abstract
3D Gaussian Splatting has emerged as an efficient photorealistic novel view synthesis method. However, its reliance on sparse Structure-from-Motion (SfM) point clouds consistently compromises the scene reconstruction quality. To address these limitations, this paper proposes a novel 3D reconstruction framework Gaussian Processes Gaussian Splatting (GP-GS), where a multi-output Gaussian Process model is developed to achieve adaptive and uncertainty-guided densification of sparse SfM point clouds. Specifically, we propose a dynamic sampling and filtering pipeline that adaptively expands the SfM point clouds by leveraging GP-based predictions to infer new candidate points from the input 2D pixels and depth maps. The pipeline utilizes uncertainty estimates to guide the pruning of high-variance predictions, ensuring geometric consistency and enabling the generation of dense point clouds. The densified point clouds provide high-quality initial 3D Gaussians to enhance reconstruction performance. Extensive experiments conducted on synthetic and real-world datasets across various scales validate the effectiveness and practicality of the proposed framework. Code is available.
@@ 1 Introduction
2 Related Work
2.1 Gaussian Processes
@@ 3 Preliminaries
4 Methodology
4.1 Overview
This paper proposes a novel framework GP-GS that enhances 3D Gaussians initialization, thereby improving 3DGS rendering quality. The overview of GP-GS is presented in Figure 2. First, addressing the limitations of sparse SfM point clouds, we propose the MOGP model to formulate point cloud densification as a regression problem. (Sec. 4.2). The training strategy of MOGP is introduced in Appendix A.1. Next, we adaptively densify sparse SfM point clouds and propose an uncertainty-based filtering strategy to remove high-uncertainty predictions. (Sec. 4.3).
4.2 Multi Output Gaussian Process
@@
where $\Gamma(\cdot)$ is the Gamma function, $\sigma^{2}$ is the variance, and $K_{\nu}(\cdot)$ is a modified Bessel function of the second kind. The parameter $\nu$ modulates how smooth the function can be: smaller $\nu$ values permit abrupt transitions, while larger $\nu$ values enforce smoother spatial changes. Details of tuning $\nu$ on various datasets can be found in Table 5. Our MOGP training procedure can be found in Appendix A.1 Algorithm 1.
4.3 Point Cloud Densification
The process of point cloud densification involves an adaptive neighbourhood-based sampling strategy, which selectively identifies pixels as candidate inputs for the MOGP model (in section 4.2). The MOGP model subsequently predicts additional point clouds and applies uncertainty filtering to enhance the density.
@@ 4.3 Point Cloud Densification
An overview of the uncertainty-based filtering procedure can be found in Appendix A.1 Algorithm 2.
5 Experiments
5.1 Dataset and Implementation Details
@@
5.2 Results of Gaussian Processes
@@
Quantitative Results To ensure the best performance of MOGP for densification, we evaluated MOGP model with different hyperparameters on our Pixel-to-Point-Cloud dataset, e.g., $\nu$ with values such as ($\nu\in\{0.5,1.5,2.5\}$). The data were split into an 80% training subset and a 20% testing subset, and the model was subsequently assessed using $R^{2}$, RMSE, and CD as we mentioned in section 5.1. The average outcomes of this analysis are summarized in Table 1, which shows that $\nu=0.5$ consistently yields the most favourable results. Consequently, we adopt $\nu=0.5$ as our default configuration for subsequent experiments. Detailed metrics across multiple datasets for this kernel selection procedure are presented in Appendix 5.2 Table 5.
@@
To validate the effectiveness of our variance-based filtering algorithm, we analyze the variance distributions across multiple Pixel-to-Point datasets, including NeRF Synthetic, Tanks & Temples, and Mip-NeRF 360, due to their diverse characteristics. The reduction in variance in all datasets suggests that our filtering method generalizes well to different types of 3D scenes, improving overall point cloud quality and enabling more accurate scene representation, as shown in Figure 5, the left column shows the spatial and colour variances before filtering on Tanks & Temples (Truck), the right column presents the variances after applying our filter. The variances in spatial components $(x,y,z)$ and colour channels $(r,g,b)$ are significantly reduced, demonstrating that our filtering process effectively removes high-variance noisy points while preserving scene consistency. More results on various datasets can be found in Appendix A.3.
@@
5.3 Quantitative and Qualitative Results on NVS
@@
Qualitative Results As shown in Figure 6, we compare our method with 3DGS* on datasets Tanks & Temples (Truck), Mip-NeRF 360 (Flowers and Bicycles), and NeRF Synthetic (Lego), we provide additional results in Appendix A.4 for each scene to further substantiate our conclusions. Our method demonstrates superior performance in NVS by effectively reconstructing fine-grained details, reducing artifacts, and enhancing occlusion handling. The visual comparisons highlight key challenges in 3DGS* rendering, including occlusions, high-frequency textures (e.g., object edges and surface details), and lighting variations. Notably, the zoom-in regions (marked with red and blue boxes) illustrate our model’s ability to recover fine structures more accurately than existing approaches. These results further validate the robustness of our method in handling complex scenes, offering improved perceptual quality and structural consistency across diverse datasets.
6 Conclusion
@@ Appendix A Appendix
A.1 MOGP Training and Dynamic Filtering Procedure
@@
A.2 MOGP Training Loss
@@
A.3 More MOGP Results
@@
A.4 More Novel View Synthesis Results